The binding energy of electrons to a chromium metal surface is 7.21 × 10-19 J. What is the longest wavelength of light (in nm) that will eject electrons from chromiummetal? What frequency is required to give electrons with kinetic energy of 5.09E-19 J? What wavelength (in nm) is required to give electrons with kinetic energy of5.09E-19 J?
Answer #1

1. Binding energy = photon energy = hc/λ

where h is Planck constant, c is speed of light and λ is wavelength

7.21 x 10-19 = 6.626 x 10-34 x 2.998 x 108

Wavelength λ = 2.76 x 10-7 m = 276 nm

2. Photon energy = binding energy + kinetic energy

= 7.21 x 10-19 + 5.09 x 10-19

= 1.23 x 10-18 J = hf

where f is frequency

Frequency f = 1.23 x 10-18/6.626 x 10-34

= 1.856 x 10151.86 x 1015 Hz

3. Wavelength λ = c/f

= 2.998 x 108/1.856 x 1015

= 1.62 x 10-7 m = 162 nm